The perimeter of a rectangular yard is 204 feet. If the width of the yard is 45 feet what is the length of the yard?
Answer by tommyt3rd(5050)

1) 2(l+w)=204
2) w=45

l+w=102
l=102-w
l=102-45
l=57 ft